How the following conversions are carried out?1. Propen to propan-1-ol 2.Ethanol to but-1-yne

Grade:12

1 Answers

vikas askiitian expert
13 years ago

1 )   propen + HBr  (in presence of peroxide)=1 bromopropan         (free radical addition)

     1bromopropane+KOH(aq)=propan-1-ol                                    (neucliophilic substitution reaction)

           0r

     propene +B2H6 +H2O2 =propan1--ol

 

2)  ethanol  + HBr(aq)= 1bromoethane   

    1bromo ethane+2Na=N-butane                (wurtz reaction)

    butane + Br2 (peroxide) =2bromo butane +KOH(alcoholic) =but-2-ene(major) + but-1-ene(minor)

    but-2-ene + Cl2 =2,3dichlorobutane

    2,3 dichlorobutane +NaNH2=but-1-yne